# On the splitting of weak nearly $\cal C$}-manifolds

### v1

@@ 4 Main results
Let $\Omega^{k}(M)$ be the space of $k$-forms on an $s$-contact manifold $(M^{2n+s},\vec{\eta},{\mathcal{R}}\oplus{\cal D})$. If $n\geq 3$, then the map $\mathcal{E}_{d\eta^{i}}:\ \Omega^{2}(M)\longrightarrow\Omega^{4}(M),\,\beta\mapsto d\eta^{i}\wedge\beta$, is injective for all $i=1,\ldots,s$.
@@
Step 1 (Symplecticity on $\mathcal{D}$ and Lefschetz injectivity): By the definition of an $s$-contact manifold, $d\eta^{i}|_{\mathcal{D}}$ is nondegenerate at every point $p\in M$. Each fibre $(\mathcal{D}_{p},d\eta^{i}|_{{\mathcal{D}}_{p}})$ is a $2n$-dimensional symplectic vector space, hence $(d\eta^{i}|_{{\mathcal{D}}})^{n}$ is nowhere vanishing. In particular, by [6, Proposition 1.1] (the algebraic Lefschetz property), the linear maps $\mathcal{E}_{d\eta^{i}|_{\mathcal{D}}}:\,\Omega^{2}(\mathcal{D})\longrightarrow\Omega^{4}(\mathcal{D})$, $\mathcal{E}_{d\eta^{i}|_{{\mathcal{D}}}}:\,\Omega^{1}(\mathcal{D})\longrightarrow\Omega^{3}(\mathcal{D})$ are injective for $n\geq 3$ and $n\geq 2$, respectively (the latter injectivity, in particular, holds when $n\geq 3$).
@@
$0=d\eta^{i}|_{\mathcal{D}}\wedge\beta$ $={d\eta^{i}|_{\mathcal{D}}\wedge\beta_{0}}+\sum\nolimits_{a=1}^{s}{\eta^{a}\wedge\bigl(d\eta^{i}|_{\mathcal{D}}\wedge\mu_{a}\bigr)}\;+\;\sum\nolimits_{1\leq a<b\leq s}{f_{ab}\,\eta^{a}\wedge\eta^{b}\wedge d\eta^{i}|_{\mathcal{D}}}.$ (73)
@@
$\bullet$ Zero $\eta$-part (first term of (73)): $d\eta^{i}|_{\mathcal{D}}\wedge\beta_{0}=0$ in $\Omega^{4}({\mathcal{D}})$. Now, as $d\eta^{i}|_{\mathcal{D}}\in\Omega^{2}({\mathcal{D}})$ and by the decomposition (72), $\beta_{0}\in\Omega^{2}({\mathcal{D}})$ and since $\mathcal{E}_{d\eta^{i}|_{\mathcal{D}}}:\Omega^{2}({\mathcal{D}})\to\Omega^{4}({\mathcal{D}})$ is injective for $n\geq 3$ (by [6, Proposition 1.1]), we get $\beta_{0}=0$.
@@
$\bullet$ Two $\eta$-parts: From Step 4 we are left with $\sum\nolimits_{1\leq a<b\leq s}f_{ab}\,\eta^{a}\wedge\eta^{b}\wedge d\eta^{i}|_{\mathcal{D}}=0$. Wedge this identity with $(d\eta^{i}|_{\mathcal{D}})^{\,n-1}$ to obtain $\sum\nolimits_{a<b}f_{ab}\,\eta^{a}\wedge\eta^{b}\wedge(d\eta^{i}|_{\mathcal{D}})^{\,n}=0$. By Step 1, $(d\eta^{i}|_{\mathcal{D}})^{\,n}$ is a nowhere-vanishing horizontal top form on $\mathcal{D}$}. Fix any point $p\in M$. Then $\{\,\eta^{a}_{p}\wedge\eta^{b}_{p}\,\}_{a<b}$ is a basis of $\Lambda^{2}(\mathrm{span}\{\eta^{1}_{p},\dots,\eta^{s}_{p}\})$, and wedging with $(d\eta^{i}|_{\mathcal{D}})^{\,n}\neq 0$ multiplies these basis elements by the same nonzero horizontal volume form. Hence, the linear independence is preserved, and the equality above forces $f_{ab}(p)=0$ for $a<b$. Since $p\in M$ is arbitrary, $f_{ab}\equiv 0$ on $M$. Thus, $\beta=\beta_{0}+\sum_{a}\eta^{a}\wedge\mu_{a}+\sum_{a<b}f_{ab}\,\eta^{a}\wedge\eta^{b}=0$, and $\mathcal{E}_{d\eta^{i}}$ is injective on $\Omega^{2}(M)$ for $n\geq 3$. ∎
The following result generalizes [18, Theorem 4.2] (see also Theorem 4.4 in [14]) to the case $s>1$.
Theorem 1.
Let $(M^{\,2n+s},{f},Q,\vec{\xi},\vec{\eta},g)$ be a weak nearly $\cal C$-manifold satisfying (5), (6), (9) and (13). If the coframe $\{\eta^{i}\}$ is uniform (has a common exterior derivative), then $n=2$ and $M$ admits a weak $f$-K-contact structure $(\hat{f},\hat{Q},\vec{\xi},\vec{\eta},g)$ defined by $\hat{f}=-\nabla\xi_{i}$ and $\hat{Q}X=R_{\xi_{i}}X\ (X\in{\cal D})$.
Proof.
$d\Phi^{(0)}_{i}=3\sum\nolimits_{j=1}^{s}\eta^{j}\wedge\Phi^{(1)}_{j},\quad d\Phi^{(1)}_{i}=3\sum\nolimits_{j=1}^{s}\eta^{j}\wedge\Phi^{(1)}_{i,j}$. (74)
Replacing $Z$ by fZ{f}Z in (49) and using (8), we obtain for all $1\leq i\leq s$
$-g((\nabla_{X}\,{f})Y,{f}^{2}h_{i}Z)=\sum\nolimits_{j=1}^{s}\big[\eta^{j}(X)\,g(h_{i}fZ,h_{j}Y)-\eta^{j}(Y)\,g(h_{j}X,{f}h_{j}Z)\big].$ (76)
By conditions, ${\bf h}\not\equiv 0$ on ${\cal D}\setminus\{0\}$, thus from (76) we get $g((\nabla_{X}\,{f})Y,V)=0$ for $X,Y,V\in{\cal D}$. By the above and (75), using $X=X^{\top}+\sum\nolimits_{i=1}^{s}\eta^{i}(X)\,\xi_{i}$ and $Y=Y^{\top}+\sum\nolimits_{i=1}^{s}\eta^{i}(Y)\,\xi_{i}$, we obtain
$g((\nabla_{X}\,{f})Y,V)$
$=\sum\nolimits_{i=1}^{s}\big[\eta^{i}(V)\,g((\nabla_{X^{\top}}\,{f})Y^{\top},\xi_{i})+\eta^{i}(X)\,g((\nabla_{\,\xi_{i}}\,{f})Y^{\top},V)+\eta^{i}(Y)\,g((\nabla_{X^{\top}}\,{f})\,\xi_{i},V)\big]$
$=\sum\nolimits_{i=1}^{s}\big[-\eta^{i}(V)\,g((\nabla_{X^{\top}}\,{f})\,\xi_{i},Y^{\top})-\eta^{i}(X)\,g((\nabla_{Y^{\top}}\,{f})\,\xi_{i},V)+\eta^{i}(Y)\,g((\nabla_{X^{\top}}\,{f})\,\xi_{i},V)\big]$
$=\sum\nolimits_{i=1}^{s}\big[\eta^{i}(V)\,g({f}h_{i}X,Y)+\eta^{i}(X)\,g({f}h_{i}Y,V)+\eta^{i}(Y)\,g({f}h_{i}V,X)\big].$ (77)
Similarly, using (4) and (45), we get
$=\sum\nolimits_{j=1}^{s}\big[\eta^{j}(X)\,g({f}\,h_{i}h_{j}Y,Z)+\eta^{j}(Y)\,g({f}\,h_{i}h_{j}Z,X)+\eta^{j}(Z)\,g({f}\,h_{i}h_{j}X,Y)\big],$
$0=d^{2}\Phi^{(0)}_{i}=3\sum\nolimits_{j=1}^{s}\big[d\eta^{j}\wedge\Phi^{(1)}_{j}-\eta^{j}\wedge d\Phi^{(1)}_{j}\big]=3\sum\nolimits_{j=1}^{s}d\eta^{j}\wedge\Phi^{(1)}_{j}$. (78)
Under the assumptions, the $\xi$-sectional curvature of $(M,g)$ is positive, which implies that the Jacobi operators $R_{\,{\xi_{i}}}:X\mapsto R_{X,\,{{\xi_{i}}}}\,{{\xi_{i}}}\ (1\leq i\leq s)$ are positive definite on $\mathcal{D}$. By [17, Theorem 2], there exists a weak $f$-K-contact structure $(\hat{f},\hat{Q},\vec{\xi},\vec{\eta},g)$ on $M$ defined by $\hat{f}=-\nabla\xi_{i}$ and $\hat{Q}X=R_{\xi_{i}}X\ (X\in{\cal D})$.
As a consequence of this, $\eta^{i}$ is a contact form and $d\eta^{i}$ is non-degenerate on $\mathcal{D}$. Thus, $M$ is a uniform $s$-contact manifold. Now, if $\dim M>4+s$, then (since $M$ is a uniform $s$-contact manifold) (78) implies that $d\eta^{j}\wedge\sum_{j=1}^{s}\Phi^{(1)}_{j}=0\,\text{for all }j.$ By Proposition 4, this condition forces $\Phi^{(1)}_{i}=0$ for all $i$; hence, ${\bf h}\equiv 0$. This contradicts the fact that $d\eta^{i}$ is non-degenerate on $\mathcal{D}$. Therefore, $\dim M\leq 4+s$. Finally, since $d\eta^{i}$ is non-degenerate on $\mathcal{D}$, the distribution $\operatorname{ker}d\eta^{i}$ has dimension $s$. We conclude that $h_{i}^{2}|_{\cal D}$ is non-zero and has at least one non-zero eigenvalue, say, $\lambda^{2}$, whose multiplicity is $4$: if $X$ is an eigenvector of $h_{i}^{2}$ with eigenvalue $\lambda^{2}$, then $\{X,fX,h_{i}X,h_{i}fX\}$ are also orthogonal eigenvectors of $h_{i}^{2}$ with the same eigenvalue. Therefore, $M$ is a $(4+s)$-dimensional manifold and the multiplicity of the eigenvalue $\lambda^{2}$ of $h_{i}^{2}$ is $4$. ∎
Remark 3.
It should be noted that the weak $f$-K-contact structure in Theorem 1 does not coincide with the existing weak nearly $\mathcal{C}$-structure $({f},Q,\vec{\xi},\vec{\eta},g)$, even though the $1$-forms $\eta^{i}$ are identical. Nevertheless, each $\eta^{i}$ is a contact form, thus $d\eta^{i}$ is non-degenerate on $\mathcal{D}$.
Corollary 2.
Let a nearly $\cal C$-manifold $M^{\,2n+s}({f},\vec{\xi},\vec{\eta},g)$ satisfy (5) and (6). If the coframe $\{\eta^{i}\}$ is uniform, then $n=2$ and $M$ admits an $f$-K-contact structure $(\hat{f},\vec{\xi},\vec{\eta},g)$ with $\hat{f}=-\nabla\xi_{i}$.
Theorem 2.
[…]
$(i)~\mathbb{R}^{s}\times\bar{M}^{\,2n},\quad(ii)~B^{\,4+s}\times\bar{M}^{\,2n-4},$ (79)
[…]
Proof.
[…]
Corollary 3.
Let $(M^{\,2n+s},{f},\vec{\xi},\vec{\eta},g)$ be a nearly $\cal C$}- (non-$\cal C$}-) manifold of $n>2$ with conditions (5) and (6). If the coframe $\{\eta^{i}\}$ is uniform, then $M$ is locally isometric (the isometry is global if $M$ is complete simply connected) to one of the Riemannian products $\mathbb{R}^{s}\times\bar{M}^{\,2n}$ and $B^{\,4+s}\times\bar{M}^{\,2n-4}$, where the induced structure on $\bar{M}$ is a nearly Kähler structure, and the induced structure on $B$ is a nearly $\cal C$}- (non-$\cal C$}-) structure.
Example 3.
Let’s illustrate (79)$(ii)$ of Theorem 2 in the case of $s=1$ (for (79)$(i)$ with $s>1$ see Example 1(iii)(iii)). Let $(f_{B},\xi_{B},\eta_{B},g_{B})$ be the nearly cosymplectic (not cosymplectic, since $d\eta_{B}\neq 0$) structure on $B^{5}=\mathbb{S}^{5}$ obtained as a totally geodesic hypersurface of the nearly Kähler six-dimensional sphere, see [4]. Define $F_{B}:=\lambda\,f_{B}$, where $\lambda\,\in\mathbb{R}_{>0}$, and $\tilde{Q}:=\lambda^{2}\,\mathrm{Id}_{B}+(1-\lambda^{2})\,\eta_{B}\otimes\xi_{B}$. Then $(F_{B},\tilde{Q},\xi_{B},\eta_{B},g_{B})$ is a weak nearly cosymplectic structure as $(\nabla^{B}_{X}F_{B})X=\lambda\,(\nabla^{B}_{X}f_{B})X=0$ for all $X\in\mathfrak{X}_{B}$, where $\nabla^{B}$ is the Levi-Civita connection on $B$. Note that, since $d\eta_{B}\neq 0$, $(F_{B},\tilde{Q},\xi_{B},\eta_{B},g_{B})$ is not weak cosymplectic. Let $(N^{2n-4},J_{N},g_{N})$ be a weak nearly Kähler manifold satisfying $J_{N}^{2}=-\lambda^{2}\operatorname{Id}_{TN}$ with $\lambda\in\mathbb{R}_{>0}$. Let $\nabla^{N}$ be the Levi-Civita connection on $N$. We equip $M=B^{5}\times N^{2n-4}$ with the product metric $g=g_{B}\oplus g_{N}$. Then the tangent bundle splits orthogonally as $TM=TB\oplus TN$, and the Levi-Civita connection on $M$ satisfies $\nabla_{X}Y=\big(\nabla^{B}_{X_{B}}Y_{B},\,\nabla^{N}_{X_{N}}Y_{N}\big)$ for $X=(X_{B},X_{N})$ and $Y=(Y_{B},Y_{N})$. The mixed derivatives on $(M,g)$ disappear since the factors are totally geodesic. For any $X=X_{B}+X_{N}\in\mathfrak{X}_{M}$, we define
$\xi:=\xi_{B},\quad\eta(X):=\eta_{B}(X_{B}),\quad f(X):=F_{B}X_{B}+J_{N}X_{N},\quad Q:=\lambda^{2}\,\mathrm{Id}+(1-\lambda^{2})\eta\otimes\xi.$
Thus $f$ acts as $F_{B}$ on the $B^{5}$-factor and as $J_{N}$ on the $N^{2n-4}$-factor; $\xi$ and $\eta$ are lifted from the $B^{5}$-factor; $\eta$ annihilates $TN^{2n-4}$: $\eta\,|_{TN}=0.$ For $X=X_{B}+X_{N}$, we have
$f^{2}(X)=F_{B}^{2}X_{B}+J_{N}^{2}X_{N}=-\lambda^{2}X_{B}-(1-\lambda^{2})\eta_{B}(X_{B})\xi_{B}+\eta_{B}(X_{B})\xi_{B}-\lambda^{2}X_{N}=-QX+\eta(X)\xi.$
Hence $f^{2}=-Q+\eta\otimes\xi$ is true, where $Q$ is a self-adjoint (1,1)-tensor. Moreover, since $g_{B}$ and $g_{N}$ satisfy their respective Hermitian identities, $g_{B}(F_{B}X_{B},F_{B}Y_{B})=g_{B}(QX_{B},Y_{B})-\eta_{B}(X_{B})\eta_{B}(Y_{B})$ and $g_{N}(J_{N}X_{N},J_{N}Y_{N})=\lambda^{2}\,g_{N}(X_{N},Y_{N})$. we have the following:
$g(fX,fY)=g(QX,Y)-\eta(X)\eta(Y),\quad\eta(\xi)=1,\quad f(\xi)=0,\quad\eta\circ f=0,\quad\eta(X)=g(X,\xi).$
This implies that $(f,Q,\xi,\eta,g)$ is a weak almost contact metric structure on $M$. Next, we verify the weak nearly cosymplectic property on $M$ as follows:
$\nabla_{X}(fX)$ $=\nabla^{B}_{X_{B}}(F_{B}X_{B})+\nabla^{N}_{X_{N}}(J_{N}X_{N}),\quad\nabla_{X}X=\nabla^{B}_{X_{B}}X_{B}{+}\nabla^{N}_{X_{N}}X_{N}.$ (80)
Applying $f$ for both sides of the last equation gives $f\nabla_{X}X=F_{B}(\nabla_{X_{B}}X_{B})+J_{N}(\nabla_{X_{N}}X_{N}).$ Combining this with (80) and the fact that $(\nabla^{B}_{X_{B}}F_{B})X_{B}=0$ and $(\nabla^{N}_{X_{N}}J_{N})X_{N}=0$, we have
$(\nabla_{X}f)X$ $=(\nabla^{B}_{X_{B}}F_{B})X_{B}+(\nabla^{N}_{X_{N}}J_{N})X_{N}=0.$
Hence, $(f,Q,\xi,\eta,g)$ is a weak nearly cosymplectic structure on $M$. Next, since $\eta$ is pulled back from $\eta_{B}$, we have $\eta=\pi_{B}^{*}(\eta_{B})$ and $d\eta=\pi_{B}^{*}(d\eta_{B})$. Since $B^{5}=\mathbb{S}^{5}$ has a weak nearly cosymplectic structure that is not weak cosymplectic, we have $d\eta_{B}\neq 0$; hence $d\eta\neq 0$. Therefore, the structure $(f,Q,\xi,\eta,g)$ is weak nearly cosymplectic but not weak cosymplectic.
Remark 4.
For $\lambda=1$, Example 3 illustrates [14, Theorem 4.5], but for $\lambda>1$, the condition (9) is not satisfied by the weak metric structure on $M$. We do not have an example of $B^{4+s}\times\bar{M}^{\,2n-4}$ with $s>1$ illustrating Theorem 2 in the case of $s>1$. Therefore, we ask the following question: Does there exist a nontrivial weak nearly $\cal C$}-manifold ($s>1$) that is the product of a weak nearly $\cal C$}-manifold $B^{\,4+s}$ and a weak nearly Kähler manifold $N^{2n-4}$?
@@ Conclusion
We have shown that the weak nearly $\cal C$}-structure is useful for studying metric $f$-structures, almost contact metric structures, and Killing vector fields. Some results on nearly cosymplectic manifolds (see [8, 14]) and weak nearly cosymplectic manifolds (see [18]) were generalized to weak nearly $\cal C$}-manifolds satisfying (9) and (13). Namely, the splitting Theorem 2 was proven and the $(4+s)$-dimensional case was characterized in Theorem 1. Their consequences (without the requirement (9) and (13)) present new results for nearly $\cal C$}-manifolds. Based on the numerous applications of the metric $f$-structure, we expect that the weak $\cal C$}-structure will also be useful for differential geometry (e.g., the theory of $\mathfrak{g}$-foliations, $s$-cosymplectic structures and $s$-contact structures) and theoretical physics.

### v2

@@ 4 Auxiliary results
Let $\Omega^{k}(M)$ be the space of $k$-forms on an $s$-contact manifold $(M^{2n+s},\vec{\eta},{\mathcal{R}}\oplus{\cal D})$. If $n\geq 3$, then the map $\mathcal{E}_{d\eta^{i}}:\ \Omega^{2}(M)\longrightarrow\Omega^{4}(M),\,\beta\mapsto d\eta^{i}\wedge\beta$ is injective for all $i=1,\ldots,s$.
@@
Step 1 (Symplecticity on $\mathcal{D}$ and Lefschetz injectivity): By the definition of an $s$-contact manifold, $d\eta^{i}|_{\mathcal{D}}$ is nondegenerate at every point $p\in M$. Each fibre $(\mathcal{D}_{p},d\eta^{i}|_{{\mathcal{D}}_{p}})$ is a $2n$-dimensional symplectic vector space, hence $(d\eta^{i}|_{{\mathcal{D}}})^{n}$ is nowhere vanishing. In particular, by [5, Proposition 1.1] (the algebraic Lefschetz property), the linear maps $\mathcal{E}_{d\eta^{i}|_{\mathcal{D}}}:\,\Omega^{2}(\mathcal{D})\longrightarrow\Omega^{4}(\mathcal{D})$, $\mathcal{E}_{d\eta^{i}|_{{\mathcal{D}}}}:\,\Omega^{1}(\mathcal{D})\longrightarrow\Omega^{3}(\mathcal{D})$ are injective for $n\geq 3$ and $n\geq 2$, respectively (the latter injectivity, in particular, holds when $n\geq 3$).
@@
$0=d\eta^{i}|_{\mathcal{D}}\wedge\beta$ $={d\eta^{i}|_{\mathcal{D}}\wedge\beta_{0}}+\sum\nolimits_{a=1}^{s}{\eta^{a}\wedge\bigl(d\eta^{i}|_{\mathcal{D}}\wedge\mu_{a}\bigr)}+\big(\sum\nolimits_{1\leq a<b\leq s}{f_{ab}\,\eta^{a}\wedge\eta^{b}\big)\wedge d\eta^{i}|_{\mathcal{D}}}.$ (78)
@@
$\bullet$ Zero $\eta$-part (the first term of (78)): $d\eta^{i}|_{\mathcal{D}}\wedge\beta_{0}=0$ in $\Omega^{4}({\mathcal{D}})$. Now, as $d\eta^{i}|_{\mathcal{D}}\in\Omega^{2}({\mathcal{D}})$ and by the decomposition (77), $\beta_{0}\in\Omega^{2}({\mathcal{D}})$ and since $\mathcal{E}_{d\eta^{i}|_{\mathcal{D}}}:\Omega^{2}({\mathcal{D}})\to\Omega^{4}({\mathcal{D}})$ is injective for $n\geq 3$ (by [5, Proposition 1.1]), we get $\beta_{0}=0$.
@@
$\bullet$ Two $\eta$-parts: From Step 4 we are left with $\big(\sum\nolimits_{1\leq a<b\leq s}f_{ab}\,\eta^{a}\wedge\eta^{b}\big)\wedge d\eta^{i}|_{\mathcal{D}}=0$. Wedge this identity with $(d\eta^{i}|_{\mathcal{D}})^{\,n-1}$ to obtain $\big(\sum\nolimits_{a<b}f_{ab}\,\eta^{a}\wedge\eta^{b}\big)\wedge(d\eta^{i}|_{\mathcal{D}})^{\,n}=0$. By Step 1, $(d\eta^{i}|_{\mathcal{D}})^{\,n}$ is a nowhere-vanishing horizontal top form on $\mathcal{D}$}. Fix any point $p\in M$. Then $\{\,\eta^{a}_{p}\wedge\eta^{b}_{p}\,\}_{a<b}$ is a basis of $\Lambda^{2}(\mathrm{span}\{\eta^{1}_{p},\dots,\eta^{s}_{p}\})$, and wedging with $(d\eta^{i}|_{\mathcal{D}})^{\,n}\neq 0$ multiplies these basis elements by the same nonzero horizontal volume form. Hence, $\{\,\eta^{a}_{p}\wedge\eta^{b}_{p}\,\wedge(d\eta^{i}|_{\mathcal{D}})^{n}\}_{a<b}$ are linearly independent $(2n+2)$-forms on $M^{2n+s}$ where $s>1$. Applying this and the equality above forces $f_{ab}(p)=0$ for $a<b$. Since $p\in M$ is arbitrary, $f_{ab}\equiv 0$ on $M$. Thus, $\beta=0$, and $\mathcal{E}_{d\eta^{i}}$ is injective on $\Omega^{2}(M)$ for $n\geq 3$. ∎
@@ Conclusion
We have shown that the weak nearly $\cal C$}-structure is useful for studying metric $f$-structures, almost contact metric structures, and Killing vector fields. Some results on nearly cosymplectic manifolds and weak nearly cosymplectic manifolds (see [10, 16, 20]) were generalized to weak nearly $\cal C$}-manifolds satisfying (10) and (13). Namely, the splitting Theorem 2 was proven and the $(4+s)$-dimensional case was characterized in Theorem 1. Their consequences (without the requirement (10) and (13)) present new results for nearly $\cal C$}-manifolds. Based on the numerous applications of the metric $f$-structure, we expect that the weak nearly $\cal C$}-structure will also be useful for differential geometry (e.g., the theory of $\mathfrak{g}$-foliations, $s$-cosymplectic structures and $s$-contact structures) and theoretical physics.